if you burn carbon with oxygen with a heat of 2050 kj how much mass will the carbon be carbon
1
Expert's answer
2017-09-21T03:24:07-0400
C + O 2 = CO 2
According to equation, 1 mole of each reactant leads to formation of 1 mole of final product.
n (C) = n (CO 2 )
Depending on the initial amount of reactants, mass of product can be calculated as:
n = m/M m = nM